# Advancing Multi-Party Dialogue Systems with Speaker-ware Contrastive Learning

### v1

@@ Abstract
Dialogue response generation has made significant progress, but most research has focused on dyadic dialogue. In contrast, multi-party dialogues involve more participants, each potentially discussing different topics, making the task more complex. Current methods often rely on graph neural networks to model dialogue context, which helps capture the structural dynamics of multi-party conversations. However, these methods are heavily dependent on intricate graph structures and dataset annotations, and they often overlook the distinct speaking styles of participants. To address these challenges, we propose CMR, a Contrastive learning-based Multi-party dialogue Response generation model. CMR uses self-supervised contrastive learning to better distinguish "who says what." Additionally, by comparing speakers within the same conversation, the model captures differences in speaking styles and thematic transitions. To the best of our knowledge, this is the first approach to apply contrastive learning in multi-party dialogue generation. Experimental results show that CMR significantly outperforms state-of-the-art models in multi-party dialogue response tasks.
Advancing Multi-Party Dialogue Systems with Speaker-ware Contrastive Learning
@@ 1 Introduction
In recent years, dialogue response generation has made notable progress Ide and Kawahara (2021); Liu et al. (2021c), largely due to advancements in language models. These improvements have enhanced the generation of coherent and contextually relevant responses in applications such as customer service, virtual assistants, and conversational agents Hu et al. (2024); Wu et al. (2022); Komeili et al. (2022). However, most research has focused on dyadic dialogues Cai and Cai (2022); Sun et al. (2021); Chen et al. (2022), involving only two participants. In contrast, multi-party dialogues, common in real-world scenarios like group discussions, meetings, and social media, present unique structural challenges. As shown in Figure 1, while dyadic dialogues follow natural speaker rotations, multi-party dialogues require careful management of information flow between multiple participants Gu et al. (2022b); Ganesh et al. (2023), posing significant challenges for response generation models.
@@
To address these complexities, many approaches utilize graph neural networks (GNNs) to model dialogue context Hu et al. (2019); Wang et al. (2020); Shi and Huang (2019). For example, Gu et al. (2022a) proposed the HeterMPC model, which uses utterance and interlocutor nodes with six meta relations to construct a heterogeneous graph that captures dialogue interactions. Similarly, Gu et al. (2023) introduced a fully connected graph to model utterance-interlocutor relationships, incorporating latent addressee inference. While GNNs effectively model the dependencies in multi-party dialogues, they have drawbacks. The complexity of graph structures and the need for extensive annotations, such as addressee labels Li and Zhao (2023), make these methods computationally expensive and challenging to scale in real-world applications.
Given these challenges, we propose CMR (Contrastive Learning-based Multi-Party Response Generation). Our approach uses contrastive learning to understand the structure and dynamics of multi-party dialogues without relying on complex graph structures or extensive annotations. The model follows a two-stage training process. In the first stage, only the encoder is trained. Through contrastive learning, the model contrasts utterances from different speakers within the same conversation and those from random speakers across the dataset. This helps the model distinguish between different facts and speaking styles, enhancing speaker awareness and leading to more relevant responses. In the second stage, the model is jointly trained on a response generation task with contrastive learning. Here, the model contrasts the dialogue context with generated responses (via beam search), the gold response, the target speaker’s previous utterance, and other speakers’ responses within the same conversation. This stage improves response quality by enhancing the model’s understanding of conversation themes.
We validated our approach using the Friends dataset. Both automatic evaluation results and human evaluation results demonstrate that our model significantly outperforms the baseline models.
@@
  • We propose CMR, a novel multi-party dialogue response generation model that leverages contrastive learning to better understand the structure and dynamics of multi-party dialogues. To the best of our knowledge, this is the first work on multi-party dialogue response generation based on contrastive learning.
  • We introduce a two-stage training process: the first stage focuses on distinguishing the speaking styles and facts among different participants through contrastive learning, and the second stage jointly trains the model with the response generation task and contrastive learning to improve the quality of responses and the model’s understanding of dialogue themes.
  • Both automatic evaluation results and human evaluation results on the Friends dataset demonstrate that our model significantly outperforms baseline models and achieves SOTA.
@@ 2.1 Graph Neural Networks for Multi-Party Dialogue
Several studies have explored the use of GNNs to capture the intricate relationships in multi-party dialogues. For instance, GSN (Graph-Structured Network) introduced by Hu et al. (2019). models multi-party dialogues as graphs rather than sequences. GSN employs an utterance-level graph-structured encoder to model information flow along the graph and incorporates speaker information to improve performance Similarly, Gu et al. (2022a) proposed the HeterMPC model, which employs heterogeneous graph networks to model interactions between interlocutors and utterances using multiple types of nodes and edges. Furthermore, another approach by Gu et al. (2023) constructed fully connected graphs to model relationships between all utterances and interlocutors in a conversation. While addressee labels are not entirely unnecessary, their model is designed to infer latent addressees through a graph prediction model integrated into a Variational Inference framework, making it robust even when some addressee information is missing.
@@ 2.2 Enhanced Speaker-Aware Approaches
Ma et al. (2023) proposed an Enhanced Speaker-Aware Multi-Party Multi-Turn Dialogue Comprehension model that integrates speaker-related features at both the token and discourse levels. By applying a speaker mask for attention and using heterogeneous graph networks for comprehensive speaker-aware discourse clues, this model achieves state-of-the-art performance on datasets such as MolweniLi et al. (2020) and FriendsQAYang and Choi (2019). Liu et al. (2021b) proposed an Utterance-Aware and Speaker-Aware Multi-turn Dialogue Representation model (MDFN) that integrates utterance-level and speaker-level features using a mask attention mechanism. The MDFN achieve amazing performance on several Datasets for response selection task. These methods highlights the importance of incorporating speaker-aware signals to improve dialogue comprehension and response generation.
@@ 2.3 Contrastive Learning for Text Generation
Contrastive learning has been effectively applied to text generation tasks to enhance the quality of generated responses. For example, the work by An et al. (2022) on contrastive neural text generation demonstrates the benefits of using contrastive objectives to distinguish between similar and dissimilar contexts, thereby improving the coherence and relevance of generated text. Additionally, Liu et al. (2021a) introduced a topic-aware contrastive learning approach for abstractive dialogue summarization. By contrasting different dialogue contexts and responses, their model better captures thematic shifts and important content, leading to more accurate summaries.
@@ 3.1 Task Formalization
Formally, let $D=\{({u_{1}},{s_{1}}),...,({u_{n}},{s_{m}})\}$ represent a multi-party conversation, which contains $n$ utterances and $m$ speakers, and $u_{i}$ is the utterance made by speaker $s_{j}$. Our objective is to develop a model $f$ that can generate a response $r$ for a specific speaker $s_{t}$ based on the context of the dialogue $D$. This can be expressed as: $r=f(D,s_{t})$. In the subsequent sections, we will describe our proposed model, CMR (Contrastive Learning-based Multi-Party Response Generation), which leverages contrastive learning in a two-stage training process to address these components effectively.
@@ 3.2 Model Overview
(a) The figure displays the training process of the stage one.
(b) The figure displays the training process of the stage two.
Figure 2: The figure displays the training processes of the CMR model. Stage One (a) shows the encoder training with contrastive learning, where the encoder processes the query and distinguishes between positive and negative samples. Stage Two (b) illustrates the joint training with response generation and contrastive learning, involving both the encoder and decoder. The combined loss of cross-entropy and InfoNCE enhances the model’s ability to generate contextually relevant responses and understand the dialogue context.
CMR is designed to effectively handle the complexities of multi-party dialogue response generation. The model employs a two-stage training process. In the first stage, the encoder is trained with contrastive learning to distinguish between the speaking styles and contexts of different participants. In the second stage, the model undergoes joint training with the response generation task and contrastive learning, enhancing its ability to generate contextually relevant and accurate responses.
3.3 Stage I: Enhancing Encoder Discrimination
The primary focus of stage I is to improve the encoder’s capacity to understand and differentiate between diverse dialogue contexts. Specifically, the encoder undergoes training using contrastive learning by contrasting the utterances of different speakers within the same conversation and those of random speakers from the entire dataset. Thus, the encoder can better distinguish between various speaking styles and related facts, enhancing its ability to capture the unique characteristics of each speaker.
3.3.1 Selection of Positive and Negative Samples
First, we randomly select an utterance from the conversation as the query $q$. The positive sample $k^{+}$ for this query is another utterance spoken by the same speaker within the same conversation. The negative samples ${k^{-}}$ consist of two parts: utterances spoken by other speakers within the same conversation, and utterances spoken by other speakers from the entire dataset, i.e., in-batch samples. It is worth noting that when selecting the query and positive and negative samples, we removed meaningless and dull replies, such as "yes yes yes," to ensure the quality and informativeness of the training data. The detailed methodology for data cleaning is provided in Appendix. By effectively distinguishing between these positive and negative samples, the encoder learns to better capture the nuances of multi-party dialogues.
3.3.2 Training Process of Stage I
The training objective for the first stage is to maximize the similarity between the query and its positive sample while minimizing the similarity between the query and its negative samples. The approach can help the model better understand "who says what" and the stylistic differences between different speakers. We achieve this by using the InfoNCE loss Oord et al. (2018) function.
@@ 3.3.2 Training Process of Stage I
$h_{q}=Pooling(Encoder(q))$ (2)
$h_{k^{+}}=Pooling(Encoder(k^{+}))$ (3)
$h_{k^{-}}=Pooling(Encoder(k^{-}))$ (4)
where $h_{q}$ is the representation of the query, $h_{k^{+}}$ is the representation of the positive sample, $h_{k^{-}_{j}}$ represents the negative samples from the same conversation, $h_{k_{i}}$ represents the samples from-batch, and $\tau$ represents temperature. Pooling refers to utilizing the hidden state of the $<$/s$>$ token to represent the encoding of the entire sequence. Because our early experiments revealed that during generation, the decoder’s cross-attention in the T5 model focuses predominantly on the $<$/s$>$ token. Consequently, the $<$/s$>$ token effectively acts as an aggregator of context information, making it a suitable representation of the entire sequence. The parameters $T$ and $N$ denote the number of negative samples from the same conversation and the batch size, respectively.
3.4 Stage II: Joint Training with Response Generation and Contrastive Learning
After the encoder has been trained to distinguish between different speakers and their styles in Stage I, the next step involves joint training with response generation and contrastive learning, which aims to further refine the model’s ability to generate contextually relevant and coherent responses.
3.4.1 Selection of Positive and Negative Samples
As with traditional contrastive learning in generation tasks, we use the dialogue context as the query and the gold response as the positive sample. However, our approach introduces an innovative method for selecting negative samples, which consists of four parts:
  • Previous Utterance by the Target Speaker: The first part of the negative samples is the utterance immediately preceding the gold response spoken by the target speaker. In general, the utterance is highly related to the gold response but contextually different enough due to the progression of the dialogue. We believe the utterance serving as a negative sample can help the model understand the dialogue context by distinguishing between closely related but semantically distinct utterances.
  • Utterances by Other Speakers in the Current Conversation: The second part includes utterances from other speakers within the same conversation with query. These samples help the model grasp the context by understanding topic shifts and interactions between different participants.
  • From Batch Samples: The third part consists of negative samples drawn from the entire dataset (batch samples). These provide a broader context and ensure that the model generalizes well across different dialogues.
  • Model-Generated Beam Search Samples: The fourth part includes the model’s own beam search samples. Following An et al. (2022), these samples, generated by the model itself, are similar in structure to the gold response but differ in content. Including these helps the model refine its understanding and improve the quality of generated responses.
3.4.2 Training Process of Stage II
In Stage Two, the model is trained to generate responses while simultaneously leveraging contrastive learning to enhance the quality of these responses. For the response generation task, we use maximum likelihood estimation with cross-entropy loss. For the contrastive learning part, we use the InfoNCE loss, similar to Stage One. Thus, the training objective in Stage Two combines the response generation loss and the contrastive learning loss. The combined loss function can be expressed as:
@@ 3.4.2 Training Process of Stage II
where $h_{q}$ is the representation of the query, $h_{k^{+}}$ is the representation of the positive sample, and $h_{k^{-}_{j}}$ represents the negative samples. The representations $h_{q}$ and $h_{k}$ for the query and the samples are obtained in a manner similar to Stage One. The parameters $P,T,N,B$ denote the number of negative samples from four different sources.
The entire training process can be referenced in Appendix 1.
@@ 4.1 Dataset
For evaluating the performance of our proposed CMR model, we considered two widely-used benchmarks in multi-party dialogue research: the Friends datasetShmueli and Ku (2019) and the Ubuntu IRC benchmarkLowe et al. (2015). The Friends dataset, which contains transcripts from the TV show "Friends," features interactions between six main characters with distinct and consistent personalities, providing a rich context for modeling speaking styles. These well-defined personas allow the model to generate dialogue that aligns with each character’s unique traits. In contrast, the Ubuntu IRC benchmark includes conversations from a large number of speakers, many of whom only contribute a single utterance. This makes it challenging to model individual speaking styles. Additionally, the Ubuntu dataset includes addressee tags and other labels that are not typically available in natural multi-party dialogues and differ from the objectives of our model. Given these factors, we determined that the Friends dataset is more suitable for evaluating our model’s ability. After data pre-processing, there are 5000 conversations for training and 653 conversations for testing.
@@ 4.2 Comparison Models
CMR 8.14 8.63 2.95 1.39 7.45
Table 1: The table shows the performance of different models on various automatic evaluation metrics. Our CMR model shows significant improvements over the baseline models. For the specific settings of the contrastive model, please refer to the appendix.
To evaluate the performance of our proposed CMR model, we compared it against several state-of-the-art models for dialogue response generation, including both general-purpose and multi-party dialogue models. Specifically, we considered BART Lewis et al. (2020) and T5 Raffel et al. (2020), both powerful transformer-based models known for their strong performance across NLP tasks, including dialogue generation. We also evaluated HeterMPC Gu et al. (2022a), which uses a heterogeneous graph neural network to capture complex interactions in multi-party dialogues, and MADNet Gu et al. (2023), which applies fully connected graph neural networks to address scenarios with missing addressee labels. Additionally, we included CONT An et al. (2022), which leverages contrastive learning to improve text generation by constructing contrastive examples from the model’s own predictions. Lastly, we used LLama3, a popular large language model with enhanced architecture and training techniques, as a baseline, testing various sizes and employing both prompt-based and in-context learning to assess its performance on our dataset.

### v2

# Advancing Multi-Party Dialogue Framework with Speaker-ware Contrastive Learning

@@ Abstract
Multi-party dialogues, common in collaborative scenarios like brainstorming sessions and negotiations, pose significant challenges due to their complexity and diverse speaker roles. Current methods often use graph neural networks to model dialogue context, capturing structural dynamics but heavily relying on annotated graph structures and overlooking individual speaking styles. To address these challenges, we propose CMR, a Contrastive learning-based Multi-party dialogue Response generation framework. CMR employs a two-stage self-supervised contrastive learning framework. First, it captures global differences in speaking styles across individuals. Then, it focuses on intra-conversation comparisons to identify thematic transitions and contextually relevant facts. To the best of our knowledge, this is the first approach that applies contrastive learning in multi-party dialogue generation. Experimental results demonstrate that CMR not only significantly outperforms state-of-the-art models, but also generalizes well to large pre-trained language models, effectively enhancing their capability in handling multi-party conversations.
Advancing Multi-Party Dialogue Framework with Speaker-ware Contrastive Learning
@@ 1 Introduction
Dialogue response generation has made notable progress Ide and Kawahara (2021); Liu et al. (2021c) in applications Hu et al. (2024); Wu et al. (2022); Komeili et al. (2022), largely due to advancements in language models. However, most research has focused on dyadic dialogues Cai and Cai (2022); Sun et al. (2021); Chen et al. (2022), involving only two participants. In contrast, multi-party dialogues, common in real-world scenarios such as group discussions, meetings, and role-play in game, present unique structural challenges. As shown in Figure 1, multi-party dialogues require careful management of information flow between multiple participants Gu et al. (2022b); Ganesh et al. (2023).
@@
To address these complexities, many approaches utilize graph neural networks (GNNs) to model dialogue context Hu et al. (2019); Wang et al. (2020); Shi and Huang (2019). For example, Gu et al. (2022a) propose the HeterMPC model, which uses utterance and interlocutor nodes with six meta relations to construct a heterogeneous graph that captures dialogue interactions. Similarly, Gu et al. (2023) introduce a fully connected graph to model utterance-interlocutor relationships, incorporating latent addressee inference. While GNNs effectively model the dependencies in multi-party dialogues, they suffer from four major limitations: (1) they require additional annotations such as speaker interaction labels; (2) they model each dialogue in isolation, failing to capture consistent speaker styles across conversations; (3) they are insensitive to topic transitions, limiting their ability to track discourse dynamics. (4) their graph-based architectures are inherently difficult to scale or integrate into large pre-trained language models.
We propose CMR (Contrastive learning-based Multi-party Response Generation), a novel framework that leverages self-supervised contrastive learning to model the structure and dynamics of multi-party dialogues without relying on complex graph structures or extensive annotations. The model follows a two-stage training process. Stage I focuses on learning to differentiate speakers based on their utterances. By contrasting utterances from different speakers within the same conversation and those from random speakers across the dataset against the target speaker’s utterances, the model learns to distinguish speaking styles among different participants. Stage II then optimizes the model for response generation with an integrated contrastive objective. Here, the model contrasts the dialogue context with generated responses (via beam search), the gold response, the target speaker’s previous utterance, and other speakers’ responses within the same conversation. This stage improves response quality by enhancing the model’s understanding of conversation themes and speaker-specific facts.
We validate our approach on both the Friends dataset and the Ubuntu Dialogue Corpus. Specifically, we evaluate CMR on two different backbone models: a traditional encoder-decoder model (T5) and a decoder-only large language model (LLaMA 3.1). Evaluation results demonstrate that our model significantly outperforms the baseline models.
@@
  • We propose CMR, a novel multi-party dialogue response generation framework that leverages contrastive learning to better understand the structure and dynamics of multi-party dialogues.
  • We introduce a two-stage training process: the first stage focuses on distinguishing the speaking styles, and the second stage improves the quality of responses and the model’s understanding of dialogue themes and speaker-specific facts.
  • Evaluation results demonstrate that our model significantly outperforms baseline models and achieves SOTA.
@@ 2.1 Graph Neural Networks for Multi-Party Dialogue
Several studies model multi-party dialogues using graph neural networks For instance, GSN (Graph-Structured Network) Hu et al. (2019) uses utterance-level graphs to capture dialogue flow and speaker dynamics. Similarly, HeterMPC Gu et al. (2022a), which employs heterogeneous graph networks to model interactions between interlocutors and utterances using multiple types of nodes and edges. Furthermore, MADNet Gu et al. (2023) utilizes fully connected graphs integrated with latent addressee inference via variational methods.
@@ 2.2 Enhanced Speaker-Aware Approaches
Ma et al. (2023) introduce a model that applies a speaker mask attention for comprehensive speaker-aware discourse clues. Liu et al. (2021b) propose an Utterance-Aware and Speaker-Aware Multi-turn Dialogue Representation model (MDFN) that integrates utterance-level and speaker-level features using a mask attention mechanism. These methods highlight the importance of incorporating speaker-aware signals.
@@ 2.3 Contrastive Learning for Text Generation
Contrastive learning has been effectively applied to text generation tasks to enhance the quality of generated responses. For example, the work by An et al. (2022) on contrastive neural text generation demonstrates the benefits of using contrastive objectives to distinguish between similar and dissimilar contexts, thereby improving the coherence and relevance of generated text. Additionally, Liu et al. (2021a) introduced a topic-aware contrastive learning approach for abstractive dialogue summarization.
@@ 3.1 Task Formalization
Formally, let $D=\{({u_{1}},{s_{1}}),...,({u_{n}},{s_{m}})\}$ represent a multi-party conversation, which contains $n$ utterances and $m$ speakers, and $u_{i}$ is the utterance made by speaker $s_{j}$. Our objective is to develop a model $f$ that can generate a response $r$ for a specific speaker $s_{t}$ based on the context of the dialogue $D$. This can be expressed as: $r=f(D,s_{t})$.
@@ 3.2 Model Overview
(a) The figure displays the training process of the Stage I.
(b) The figure displays the training process of the Stage II.
Figure 2: The figure displays the training processes of the CMR framework. Here, an encoder-decoder architecture is presented as an example. Stage I (a) shows the encoder training with contrastive learning. Stage II (b) illustrates the joint training with response generation and contrastive learning.
The framework employs a two-stage training process. In the first stage, the model learns utterance-level representations using contrastive learning, enabling it to distinguish between the speaking styles and contexts of different participants. In the second stage, the model undergoes joint training with the response generation task and contrastive learning, enhancing its understanding of dialogue themes and speaker-specific facts to generate contextually relevant and accurate responses.
To provide a detailed understanding of the model’s mechanics, we divide the discussion into two sections: Stage I (Sec 3.3) and Stage II(Sec 3.4). Each section covers the selection of positive and negative samples, as well as the training process specific to that stage.
3.3 Stage I: Speaker Discrimination Learning
We use contrastive learning to differentiate speaker styles, contrasting utterances within and across conversation.
3.3.1 Stage I: Selection of Samples
We select samples for contrastive learning as follows:
  1. 1.
    Query ($q$): An utterance randomly selected from a conversation.
  2. 2.
    Positive Sample ($k^{+}$): Another utterance spoken by the same speaker within the same conversation.
  3. 3.
    Negative Samples (${k^{-}}$):
    • Utterances spoken by other speakers within the same conversation.
    • Utterances from random speakers across the dataset (in-batch samples).
It is worth noting that when selecting the query and positive and negative samples, we remove meaningless and dull replies, such as "yes yes yes," to ensure the quality and informativeness of the training data. The detailed methodology for data cleaning is provided in Appendix B.
3.3.2 Stage I: Training Objective and Process
We optimize the InfoNCE loss Oord et al. (2018) to bring utterances from the same speaker closer while pushing apart those from different speakers, encouraging the model to better understand "who says what" and stylistic differences between different speakers:
@@ 3.3.2 Stage I: Training Objective and Process
where the parameters $T$ and $N$ denote the number of negative samples from the same conversation and the batch size, respectively. $h_{q}$ represents the embedding of the query utterance, $h_{k^{+}}$ is the embedding of the positive sample, and $h_{k^{-}_{j}}$ are embeddings of negative samples drawn from the same conversation, $h_{k_{i}}$ represents the in-batch samples, and $\tau$ represents the temperature.
Specifically, for encoder-decoder architectures, these embeddings are obtained by encoding each utterance separately through the encoder and applying a pooling operation over the encoder’s token-level hidden states. Thus, only encoder are trained for encoder-decoder model in Stage I. For decoder-only architectures, the embeddings are obtained by passing each utterance through the transformer-decoder individually and applying mean-pooling across the token-level hidden states produced by the decoder.
3.4 Stage II: Enhancing Contextual Response
This stage involves joint training with response generation and contrastive learning, which aims to further refine the model’s ability to generate contextually relevant and coherent responses.
3.4.1 Stage II: Selection of Samples
For a specific conversation, we select samples of the following types:
  1. 1.
    Query ($q$): Dialogue context.
  2. 2.
    Positive Sample ($k^{+}$): The gold response for the target speaker.
  3. 3.
    Negative Samples ($k^{-}$): The negative samples combine those from Stage I with additional types:
    • The target speaker’s previous utterance.
    • Model-generated beam search samples.
3.4.2 Stage II: Training Objective and Process
For the response generation task, we use maximum likelihood estimation with cross-entropy loss. For the contrastive learning part, we use the InfoNCE loss, similar to Stage I. Thus, the training objective in Stage II combines the response generation loss and the contrastive learning loss. The combined loss function can be expressed as:
@@ 3.4.2 Stage II: Training Objective and Process
The representations $h_{q}$ and $h_{k}$ for the query and samples are obtained in a manner similar to Stage I. The parameters $P,T,N,B$ denote the number of negative samples from four different sources.
The entire training process can be referenced in Appendix A.
@@ 4.1 Dataset
To evaluate the performance of CMR, we consider two widely-used benchmarks in multi-party dialogue research: the Friends datasetShmueli and Ku (2019) and the Ubuntu IRC benchmarkLowe et al. (2015). The Friends dataset, which contains transcripts from the TV show "Friends," features interactions between six main characters with distinct and consistent personalities, providing a rich context for modeling speaking styles. In contrast, the Ubuntu dataset includes many speakers with sparse utterances, posing challenges for style modeling. Given these factors, we determine that the Friends dataset is more suitable for evaluating our model’s ability. However, to assess the generalization capability of our model, we apply the second-stage trained model to the Ubuntu dataset.
@@ 4.2 Comparison Models
LLaMA3-70b-lora 5.33 5.12 2.11 0.98 4.90
CMR-T5 8.14 8.63 2.95 1.39 7.45
CMR-LLaMA3 8b 8.71 9.13 3.16 1.44 7.82
Ubuntu
GSN*Hu et al. (2019) - 9.02 2.78 - -
ASRG*Song et al. (2022) - 9.41 2.89 - -
ChatMDG*Li et al. (2024) - 9.72 3.04 - -
BartLewis et al. (2020) 11.71 11.07 3.75 1.62 10.35
T5Raffel et al. (2020) 12.04 11.91 3.65 1.74 10.69
HeterMPCGu et al. (2022a) 11.95 11.58 3.87 1.71 10.46
MADNetGu et al. (2023) 12.13 11.64 3.81 1.78 11.13
LLaMA3-8b-lora 12.75 12.39 3.98 1.80 12.07
LLaMA3-70b-lora 12.61 12.43 4.03 1.66 12.00
CMR-T5(only Stage II) 12.41 12.37 3.99 1.80 11.89
CMR-LLaMA3 8b (only Stage II) 13.43 13.01 4.28 1.86 12.89
Table 1: The table shows the performance of different models on various automatic evaluation metrics. For the specific settings of the LLaMA3 model, please refer to the appendix C. * indicates that we directly report the results obtained from training on the original UBUNTU dataset as reported in the original paper.
To evaluate the performance of CMR, we compare it against several state-of-the-art models for dialogue response generation, including both general-purpose and multi-party dialogue models. Evaluated models include graph-based methods (GSNHu et al. (2019), HeterMPC Gu et al. (2022a), MADNet Gu et al. (2023), ChatMDGLi et al. (2024)), contrastive learning methods (CONT An et al. (2022)), and transformer-based models (BART Lewis et al. (2020), T5 Raffel et al. (2020)). Notably, BART and T5 also serve as the backbone for all the aforementioned architectures. Lastly, we evaluate LLaMA3.1, a popular large language model.